Solve the given differential equation by using an appropriate substitution. The DE is of the form dy dx = f(Ax + By + C)




dy/dx = cos(x + y)

1
Expert's answer
2021-11-11T12:23:52-0500

"u=x+y"

"y'=u'-1"

"u'-cosu-1=0"


"\frac{u'}{cosu+1}=1"


"\frac{du}{cosu+1}=dx"


"\int\frac{du}{cosu+1}=\int dx"


"tan(u\/2)=x+c"


"tan((x+y)\/2)=x+c"